OpenStudy (anonymous):

prove: cos^2x/sinx+sinx=cscx

OpenStudy (anonymous):

multiply all terms by sinx

OpenStudy (anonymous):

\[\cos^2x+\sin^2x=cscxsinx\] \[\cos^2x=\sin^2x=1\]

OpenStudy (anonymous):

\[\cos^2x+\sin^2x=1\] Typo above, this is a pythagorean identity

OpenStudy (anonymous):

so that is overall answer

OpenStudy (anonymous):

well this is an identity

OpenStudy (anonymous):

i know it is asking me to prove the identity

OpenStudy (anonymous):

ok